(a+b)^2=2ab+c^2 is the problem. I am told to subtract 2ab from each side. What is the final answer?
Thanks,
tate

Answer by jim_thompson5910(35256)

(a+b)^2=2ab+c^2


(a+b)^2-2ab=c^2


a^2+2ab+b^2-2ab=c^2


a^2+b^2 = c^2